OpenStudy (anonymous):

what did i do wrong here The altitude of a triangle is increasing at a rate of 1.5 centimeters/minute while the area of the triangle is increasing at a rate of 4.5 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 7 centimeters and the area is 90 square centimeters? Area = A, altiitude = h and base = b => A = (1/2)bh => dA/dt = (1/2) (b*dh/dt + h*db/dt) ... ( 1 ) A = 90 sq. cm and h = 7 cm => 90 = (1/2) b*7 => b = 180/7 Plugging, dA/dt = 4.5, dh/dt = 1.5, h = 7 and b = 180/7 in eqn. ( 1 ), 4.5 = (1/2) [(180/7) 1.5 + 7 db/dt) => 9

OpenStudy (anonymous):

=> 9 = (270/7) + 7 db/dt => db/dt = (1/7) (9 - 270/7) = - 207/49 cm/minute => base is decreasing at the rate of 207/49 cm/minute.

OpenStudy (anonymous):

Area = A, altiitude = h and base = b => A = (1/2)bh => dA/dt = (1/2) (b*dh/dt + h*db/dt) ... ( 1 ) A = 90 sq. cm and h = 7 cm => 90 = (1/2) b*7 => b = 180/7 Plugging, dA/dt = 4.5, dh/dt = 1.5, h = 7 and b = 180/7 in eqn. ( 1 ), 4.5 = (1/2) [(180/7) 1.5 + 7 db/dt) => 9 = (270/7) + 7 db/dt => db/dt = (1/7) (9 - 270/7) = - 207/49 cm/minute => base is decreasing at the rate of 207/49 cm/minute.

OpenStudy (bahrom7893):

did u just copy pate? wth?

OpenStudy (bahrom7893):

*paste

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

What are you saying yo

OpenStudy (anonymous):

the final ans in that solution is wrong...whats the cause?

OpenStudy (anonymous):

altitude-rate = 1.5 cm/min area-rate = 4.5 cm^2/min altitude = 7 cm area = 90 cm^2 Let, base = b cm, SO, (1/2)(b)(7) = 90, b = 90 * 2 / 7 = 180/7 cm Now, altitude-time, t1 = 7/1.5 = 14/3 min area-time, ta = 90/4.5 = 20 min Let, base-time = tb, SO, (1/2)(tb)(t1) = ta, SO, tb = 20*2 / (14/3) = 60/7 min Hence, base-rate = base/base-time = (180/7) / (60/7), base-rate = 3 cm/min >==================< ANSWER

OpenStudy (bahrom7893):

okay... uhmm lets see...

OpenStudy (bahrom7893):

=> 9 = (270/7) + 7 db/dt wrong..

OpenStudy (bahrom7893):

you had a half already... remember? but u're multiplying the inside out again.. just drop the half and that's it

OpenStudy (bahrom7893):

wait nevermind.. i was dumb

OpenStudy (anonymous):

A = (1/2) * h * b Chain rule dA/dt = (1/2) * dh/dt * b + (1/2) * h * db/dt We know the following dA/dt = 4.5 cm^2/min dh/dt = 1.5 cm/min h = 7 cm b = 2A/h = 2 * 90 cm^2 / 7 cm = 25 5/7 cm dA/dt = (1/2) * dh/dt * b + (1/2) * h * db/dt 4.5 cm^2/min = (1/2) * 1.5 cm/min * 25 5/7 cm + (1/2) * 7 cm * db/dt 9 cm^2/min = 38 4/7 cm^2/min + 7 cm * db/dt 9 cm^2/min - 38 4/7 cm^2/min = 7 cm * db/dt -29 4/7 cm^2/min = 7 cm * db/dt -29 4/7 cm^2/min / 7 cm = db/dt -5 25/49 cm/min = db/dt

OpenStudy (bahrom7893):

= (1/7) (9 - 270/7) != - 207/49 cm/minute it is = -261/49 cm/min

OpenStudy (anonymous):

hmm, webwork says thats not right either

OpenStudy (anonymous):

Bahrom what are you doing

OpenStudy (anonymous):

kumar what do you mean

OpenStudy (bahrom7893):

ok u guys finally forced me to write this out...

OpenStudy (anonymous):

haha sorry bahrom

OpenStudy (anonymous):

Sorry for what?

OpenStudy (anonymous):

sorry for forcing him to write it out

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

whats the wow for?

OpenStudy (anonymous):

Wow for u saying sorry what else can a wow be fore.

OpenStudy (anonymous):

hes just voluntarily helping me out. hes not getting paid so...

OpenStudy (anonymous):

Oh okay

OpenStudy (bahrom7893):

jinnie, your original answer was correct. all modifications were wrong. webworks sucks, tell the professor to go and commit a suicide for assigning these problems

OpenStudy (anonymous):

Bahrom it's fine don't worry about it

OpenStudy (anonymous):

man this sucks. this thing is due tomorrow. I cant meet the professor before then. f*******